Fermi National Laboratory
 

Operation Summary for January 14, 2011 - January 17, 2011

 

The Integrated Luminosity for the period from 1/10/11 to 1/17/11 was 64.13 pb-1 (inverse picobarns). NuMI reported receiving 7.82E18 protons on target during this same period.

Friday Jan 14

The day shift began with Operations monitoring the Tevatron (TEV) store 8419 (227.64E30) and the Recycler stash (65.8E10), with the Antiproton Source (Pbar) stacking (20.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, at 10 AM, Operations halted beam to Pbar and NuMI for chilled water system repairs at the Central Utilities Building (CUB). Operations terminated store 8419 at 10:50 AM. Operations reported that a power supply over temperature trip shut off beam to T-978.

During the evening shift, Operations monitored the accelerator complex. MiniBooNE was the only experiment taking beam.

Saturday Jan 15

The midnight shift began with the CUB chilled water system work in progress, with Operations monitoring the stash (109.3E10) and the Pbar stack (10E10), with Pbar and NuMI off due to CUB work, and with MiniBooNE taking beam.

During the midnight shift, FESS personnel reported at 1:45 AM that they had completed the CUB chilled water system repairs. Operations restored beam to Pbar and NuMI at 3:30 AM. Operations established store 8423 at 6:30 AM with an average initial luminosity (AIL) of 176.75E30.

During the day shift, at 12:54 PM, Operations reported that store 8423 aborted due to the failure of a TEV correction element power supply. EE Support personnel repaired the power supply. Operations established store 8425 at 4:53 PM with an AIL of 241.8E30.

During the evening shift, Operations monitored the accelerator complex.

Sunday Jan 16

The midnight shift began with Operations monitoring store 8425 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operations terminated store 8425 at 8:34 AM. Operations established store 8426 at 9:43 AM with an AIL of 346.3E30.

During the evening shift, Operations monitored the accelerator complex.

Monday Jan 17

The midnight shift began with Operations monitoring store 8426 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8426 at 3:25 AM. Operations established store 8427 at 4:33 AM with an AIL of 336.11E30.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/18/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory